# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ 5.1.1 Linear Layers Mimic Deep Layers in MLPs for Feature Learning
In this subsection, we study the roles of linear layers and MLP layers at deep layers in feature learning. To begin, we provide the experimental setup for our experiments.
@@
In these experiments, we construct two 7-layer networks: (a) a nonlinear MLP network and (b) a hybrid network formed by a 3-layer MLP followed by a 4-layer linear network. We set the hidden dimension $d=1024$ for all linear layers.
@@
We employ the SGD optimizer to train the networks by minimizing the MSE loss on the CIFAR-10 dataset [67]. For the settings of the SGD optimizer, we use a momentum of 0.9, a weight decay of $10^{-4}$, and a dynamically adaptive learning rate ranging from $10^{-3}$ to $10^{-5}$, modulated by a CosineAnnealing learning rate scheduler as detailed in [76]. We use the orthogonal initialization in (18) with $\xi=0.1$ to initialize the network weights. The neural networks are trained for 200 epochs with a batch size of 128.
@@
Now, we elaborate on the tasks conducted in Figures 1, 2, 5, 7, and 8 and draw conclusions from our observations.
  • In Figure 1 (resp. Figure 7), we plot the numerical rank and training accuracy of the features against the layer index after training (resp. before training). Here, the numerical rank of a matrix is defined as the number of its singular values whose sum collectively accounts for more than 95% of its nuclear norm. In terms of training accuracy, we add a linear classification layer to a given layer $l$ of the neural network and train this added layer using the cross-entropy loss to compute the classification accuracy. It is observed in Figure 1 that the training accuracy rapidly increases in the initial layers and nearly saturates in the deep layers, whereas the numerical rank increases in the initial layers and then decreases in the deep layers progressively in the trained networks. This observation suggests that the initial layers of a network create linearly separable features that can achieve accurate classification, while the subsequent layers further compress these features progressively.
  • In Figure 2, we employ a 2D UMAP plot [81] with the default settings in the UMAP Python package to visualize the evolution of features from shallow to deep layers. It is observed that in the first layer, features do not exhibit obvious structures in terms of feature compression and discrimination; the features from the same class are more and more compressed, while the features from different classes become more and more separable from layer 2 to layer 4; this within-class compression and between-class discrimination pattern is strengthened from layer 4 to layer 6. This visualization demonstrates that features in the same class are compressed, while the features from different classes are discriminated progressively.
  • In Figure 5 (resp. Figure 8), we plot the metrics of within-class feature compression $C_{l}$ and between-class feature discrimination $D_{l}$ defined in 1 against the layer index after training (resp. before training). According to Figure 5, we observe a consistent trend of decreasing $C_{l}$ and increasing $D_{l}$ against the layer index after training for both networks, albeit the hybrid network exhibits a smoother transition in $C_{l}$ and $D_{l}$ compared to the fully nonlinear network. This observation supports our result in 1.
@@
Comparing Figure 1(a) to Figure 1(b) and Figure 5(a) to Figure 5(b), we conlcude that linear players play the same role as MLP layers in the deep layers of nonlinear networks in feature learning, compressing within-class features and discriminating between-class features progressively. Intuitively, this is because the representations from the initial layers are already linearly separable, we can replace the deeper nonlinear layers with DLNs to achieve the same functionality without sacrificing the training performance. Moreover, comparing the results after training (Figures 1 and 5) with those before training (Figures 7 and 8), we conclude that training neural networks using SGD learn feature representations beyond NTK regime. This finding aligns with the results in recent works [12, 32].
@@ 5.1.2 Effects of Depth in Linear Layers for Improving Generalization
In this subsection, we study the impact of depth of deep networks on generalization performance. To begin, we provide the experimental setup for our experiments.
@@
We consider a 2-layer MLP network as our base network and construct networks of varying depths from 3 to 9 by adding linear or MLP layers to this base network. We use the PyTorch default initialization [57] to initiate all weights.
@@
We train these networks on the FashionMNIST [120] and CIFAR-10 [67] datasets, using the same training approach in Section 5.1.1, except that the initial learning rate is set as $10^{-2}$. We use $5$ different random seeds to initialize the network weights and report the average test accuracy.
@@
After training, we report the average test accuracy throughout training against the number of layers of constructed networks in Figure 3. Notably, we observe a consistent improvement in test accuracy as the networks grow deeper, regardless of whether the added layers are linear or MLP layers. This observation provides further evidence for the resemblance between DLNs and the deeper layers in MLPs. Additionally, it also shows that DLNs can benefit from increased depth, similar to their counterparts in traditional nonlinear networks.
5.2 Experimental Verification of 2 and 1
@@
In all experiments, we fix the number of classes $K=3$ and the number of training samples $N=30$, with varying hidden dimensions $d\geq N$ as specified in each figure. To generate input data $\bm{X}\in\mathbb{R}^{d\times N}$ that satisfies Assumption 1, we first generate a matrix $\bm{A}\in\mathbb{R}^{d\times N}$ with entries i.i.d. sampled from the standard normal distribution and a matrix $\bm{B}\in\mathbb{R}^{d\times N}$ with entries i.i.d. sampled from the uniform distribution on $[0,1]$. Then, we apply a compact SVD to $\bm{A}$ and set ${\bm{X}}_{0}\in\mathbb{R}^{d\times N}$ as the left singular matrix. Next, we obtain a matrix $\bm{N}$ by normalizing $\bm{B}$ such that its Frobenious norm is 1. Finally, we generate $\bm{X}$ via $\bm{X}={\bm{X}}_{0}+\bm{N}$ such that $\bm{X}$ is nearly orthogonal satisfying 1.
@@
5.2.1 Experiments for Verifying 2
In this subsection, we conduct experiments to verify 2 and corroborate the discussions in Section 3.2 based on the above experimental setup. Specifically, we train DLNs using orthogonal initialization in (18) with $\xi=0.1$ for various network depths and widths. Given a DLN, we train it 10 times with different random seeds. In each run, we terminate training once the training loss is less than $10^{-11}$. We plot the following metrics over 10 runs against depth and width in Figure 6(a) and Figure 6(b), respectively:
[…]
where $\mu=\frac{1}{(L-1)(d-2K)}\sum_{l=1}^{L-1}\sum_{i=K+1}^{d-K}\sigma_{i}(\bm{W}_{l})$. It can be observed from Figure 6 that the weights of a DLN with different depths and widths trained under our settings approximately satisfy 2. This supports our discussions in Section 3.2.
5.2.2 Experiments for Verifying 1
Now, we conduct experiments to validate 1. Specifically, we train 6-layer DLNs using GD with varying initialization scales $\xi=0.1,0.3,0.5$, respectively. Then, we plot the metrics of within-class compression $C_{l}$ and between-class discrimination $D_{l}$ against layer index in Figure 9(a) and Figure 9(b), respectively. It is observed that the feature compression metric $C_{l}$ decreases exponentially, while the feature discrimination metric $D_{l}$ increases linearly w.r.t. the layer index. In the top row of Figure 9, the solid blue line is plotted by fitting the values of $C_{l}$ at different layers, and the green and orange dash-dotted lines are plotted according to the lower and upper bounds in (15), respectively. We can observe that the solid blue line is tightly sandwiched between the green and orange dash-dotted lines. This indicates that (14) and (15) provide a valid and tight bound on the decay rate of the feature compression metric. According to the bottom row of Figure 9, we conclude that (16) provides a valid bound on the growth rate of the feature discrimination metric.
Moreover, we respectively train 9-layer DLN and MLP networks with hidden dimension $d=50$ via GD with an initialization scale $\xi=0.3$. In Figure 4, we plot the feature compression and discrimination metrics on both DLN and MLP networks, respectively. We can observe that exponential decay of feature compression and linear increasing of feature discrimination holds exactly on DLNs, and approximately on nonlinear networks.
@@ 5.3.1 Empirical Results Beyond Theory
  • Progressive feature compression and discrimination in nonlinear deep networks. We train 8-layer and 16-layer MLP networks, respectively. After training, we plot the metrics of feature compression and discrimination defined in (1) in Figure 10. It is observed that the feature compression metric $C_{l}$ decays at an approximate geometric rate in nonlinear networks, while the feature discrimination metric $D_{l}$ increases at an approximate linear rate. Additionally, it is worth mentioning that a similar “law of separation” phenomenon has been reported in [53, 79] on nonlinear networks, but their results are based upon a different metric of data separation.
@@
Beyond these findings, we conduct additional experiments on the SST-5 text dataset [102] and the ImageNet [28] dataset to further support our claims. These results are reported in Section C.2.
(a) 8-layer MLP.
(b) 16-layer MLP.
@@
(a) 5-layer DLN with Kaiming Init.
(b) 8-layer DLN with Kaiming Init.
@@
(a) FashionMNIST
(b) CIFAR-10
Figure 12: Progressive feature compression on DLNs trained with default initialization and real datasets. Using the DLNs trained in Figure 3, we plot the within-class compression metrics $C_{l}$ against the layer indices. It is observed that progressive linear decay still (approximately) happens without the orthogonal initialization and datasets described in 1.
@@ 5.3.2 Implication on Transfer Learning
In this subsection, we experimentally substantiate our claims on transfer learning in Section 3 on practical nonlinear networks and real datasets, demonstrating that features before projection heads are less collapsed and exhibit better transferability. To support our argument, we conduct our experiments based on the following setup.
@@
We employ a ResNet18 backbone architecture [58], incorporating $t\in\{1,2,3,4,5\}$ layers of projection heads between the feature extractor and the final classifier, respectively. Here, one layer of the projection head consists of a linear layer followed by a ReLU activation layer. These projection layers are only used in the pre-training phase. On the downstream tasks, they are discarded, and a new linear classifier is trained on the downstream dataset.
@@
We use the CIFAR-100 and CIFAR-10 dataset in the pre-training and fine-tuning tasks, respectively. We train the networks using the Rescaled-MSE loss [47], with hyperparameters set to $k=5$ and $M=20$ for 200 epochs. During pre-training, we employed the SGD optimizer with a momentum of 0.9, a weight decay of $5\times 10^{-4}$, and a dynamically adaptive learning rate ranging from $10^{-2}$ to $10^{-5}$, modulated by a CosineAnnealing learning rate scheduler [76]. During the fine-tuning phase, we freeze all the parameters of the pre-trained model and only conduct linear probing. In other words, we only train a linear classifier on the downstream data for an additional 200 epochs. We run each experiment with 3 different random seeds.
@@
Figure 13: Adding projection heads mitigates feature collapse and improves transfer accuracy. We pre-train ResNet18 backbones with 1 to 5 layers of projection heads on the CIFAR-100 dataset. After pre-training, we drop the projection heads and compute the metric $C_{l}$ on the learned features. Finally, we linearly probe the pre-trained models on the CIFAR-10 dataset and plot the evolution of $C_{l}$ alongside the corresponding transfer learning performance against the number of projection head layers. The shaded area represents the standard deviation of $3$ random seeds. The plot shows a clear trend that adding more layers of projection heads mitigates feature collapse and improves transfer learning results.
@@ 6 Conclusion
In this work, we studied hierarchical representations of deep networks by investigating intermediate features. In the context of training DLNs for solving multi-class classification problems, we explored the evolution of features across layers by analyzing within-class feature compression and between-class feature discrimination. We showed that under mild assumptions regarding the input data and trained weights, each layer of a DLN progressively compresses within-class features at a geometric rate and discriminates between-class features at a linear rate w.r.t. the layer index. Moreover, we discussed the implications of our results for deep learning practices related to interpretation, architecture design, and transfer learning. Our extensive experimental results on synthetic and real data sets not only support our theoretical findings but also illuminate our results regarding deep nonlinear networks.
@@
Our work has opened many interesting directions to be further explored. First, our extensive experiments have demonstrated analogous patterns of progressive compression and discrimination phenomena in deep nonlinear networks. Extending our analysis to deep nonlinear networks emerges as a natural and promising direction for future research. Second, [121, 45] demonstrated that the outputs of self-attention layers in transformers exhibit a similar progressive compression phenomenon. It would be interesting to study this phenomenon in transformers based on our proposed framework.
Acknowledgment
@@
XL, CY, PW and QQ acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, an AWS AI Award, a gift grant from KLA, and MICDE Catalyst Grant. PW and LB acknowledge support from DoE award DE-SC0022186, ARO YIP W911NF1910027, and NSF CAREER CCF-1845076. ZZ acknowledges support from NSF grants CCF-2240708 and IIS-2312840. WH acknowledges support from the Google Research Scholar Program. Results presented in this paper were obtained using CloudBank, which is supported by the NSF under Award #1925001. The authors acknowledge valuable discussions with Dr. Weijie Su (Upenn), Dr. John Wright (Columbia), Dr. Rene Vidal (Upenn), Dr. Hangfeng He (U. Rochester), Mr. Zekai Zhang (Tsinghua U.), Mr. Yaodong Yu (UC Berkeley), and Mr. Yuexiang Zhai (UC Berkeley) at various stages of the work.
Supplementary Material
Appendix A Auxiliary Results
Lemma 1.
Proof.
Let $\bm{A}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decomposition of $\bm{A}$, where
$\bm{U}\in\mathcal{O}^{m}$, and $\bm{V}\in\mathcal{O}^{n}$. Then we compute
which, together with the AM-QM inequality, directly implies (21). ∎
Lemma 2.
Lemma 3.
$\bm{Q}\bm{Q}^{T}=\bm{U}\bm{U}^{T}.$ (23)
Proof.
It follows from $\bm{U}\bm{\Lambda}\bm{U}^{T}=\bm{Q}\bm{A}\bm{Q}^{T}$, $\bm{Q}\in\mathcal{O}^{d\times K}$, and $\bm{U}\in\mathcal{O}^{d\times K}$ that
$\bm{Q}^{T}\bm{U}\bm{\Lambda}=\bm{A}\bm{Q}^{T}\bm{U}.$ (24)
Now, left-multiplying both sides by $\bm{Q}$ yields
Lemma 4.
For every real number $1\leq\theta\leq 1$ and $x\geq-1$, it holds that
$(1+x)^{\theta}\leq 1+\theta x.$
Appendix B Proofs of Main Results

### v4

@@ 5.1.1 Linear Layers Mimic Deep Layers in MLPs for Feature Learning
In this subsection, we study the roles of linear layers and MLP layers in deep layers for feature learning. To begin, we provide the experimental setup for our experiments.
@@
In these experiments, we construct two 8-layer networks: (a) a nonlinear MLP network and (b) a hybrid network formed by a 3-layer MLP followed by a 5-layer linear network. In both cases, the final layer ($L$-th layer) serves as a linear classifier. We set the hidden dimension $d=1024$ for all linear layers.
@@
We employ the SGD optimizer to train the networks by minimizing the MSE loss on the CIFAR-10 dataset (krizhevsky2009learning). For the settings of the SGD optimizer, we use a momentum of 0.9, a weight decay of $10^{-4}$, and a dynamically adaptive learning rate ranging from $10^{-3}$ to $10^{-5}$, modulated by a CosineAnnealing learning rate scheduler as detailed in loshchilov2017sgdr. We use the orthogonal initialization in (18) with $\xi=0.1$ to initialize the network weights. The neural networks are trained for 400 epochs with a batch size of 128.
@@
Now, we elaborate on the tasks conducted in Figures 1, 2, 6, 8, and 9 and draw conclusions from our observations.
  • In Figure 1 (resp. Figure 8), we plot the numerical rank and training accuracy of the features against the layer index after training (resp. before training). Here, the numerical rank of a matrix is defined as the number of top singular values whose sum collectively accounts for more than 95% of its nuclear norm. In terms of training accuracy, we add a linear classification layer to a given layer $l$ of the neural network and train this added layer using the cross-entropy loss to compute the classification accuracy. It is observed in Figure 1 that the training accuracy rapidly increases in the initial layers and nearly saturates in the deep layers, whereas the numerical rank increases in the initial layers and then decreases in the deep layers progressively in the trained networks. This observation suggests that the initial layers of a network create linearly separable features that can achieve accurate classification, while the subsequent layers further compress these features progressively.
  • In Figure 2, we employ a 2D UMAP plot (mcinnes2018umap) with the default settings in the UMAP Python package to visualize the evolution of features from shallow to deep layers. It is observed that in the first layer, features do not exhibit obvious structures in terms of feature compression and discrimination; the features from the same class are more and more compressed, while the features from different classes become more and more separable from layer 2 to layer 4; this within-class compression and between-class discrimination pattern is strengthened from layer 4 to layer 6. This visualization demonstrates that features in the same class are compressed, while the features from different classes are discriminated progressively.
  • In Figure 6 (resp. Figure 9), we plot the metrics of within-class feature compression $C_{l}$ and between-class feature discrimination $D_{l}$ defined in Definition 1 against the layer index after training (resp. before training). According to Figure 6, we observe a consistent trend of decreasing $C_{l}$ and increasing $D_{l}$ against the layer index after training for both networks, albeit the hybrid network exhibits a smoother transition in $C_{l}$ and $D_{l}$ compared to the fully nonlinear network. This observation supports our result in Theorem 4.
@@
Comparing Figure 1(a) to Figure 1(b) and Figure 6(a) to Figure 6(b), we conclude that linear layers play the same role as MLP layers in the deep layers of nonlinear networks in feature learning, compressing within-class features and discriminating between-class features progressively. Intuitively, this is because the representations from the initial layers are already linearly separable, hence we can replace the deeper nonlinear layers with DLNs to achieve the same functionality without sacrificing the training performance. Moreover, comparing the results after training (Figures 1 and 6) with those before training (Figures 8 and 9), we conclude that neural networks trained with SGD do not operate in the lazy-training regime described by the neural tangent kernel (NTK) theory, where network parameters remain close to their initialization and training behaves nearly linearly. This observation is consistent with findings from recent studies (bietti2022learning; Damian2022).
@@ 5.1.2 Effects of Depth in Linear Layers for Improving Generalization
In this subsection, we study the impact of depth in deep networks for generalization. To begin, we provide the experimental setup for our experiments.
@@
We consider a 2-layer MLP network as our base network and construct networks of varying depths from 3 to 9 by adding linear or MLP layers to this base network. We use the PyTorch default initialization (he2015delving) to initiate all weights.
@@
We train these networks on the FashionMNIST (xiao2017fashion) and CIFAR-10 (krizhevsky2009learning) datasets, using the same training approach in Section 5.1.1, except that the initial learning rate is set as $10^{-2}$. We use $5$ different random seeds to initialize the network weights and report the average test accuracy.
@@
After training, we report the average test accuracy throughout training against the number of layers of constructed networks in Figure 3. Notably, we observe a consistent improvement in test accuracy as the networks grow deeper, regardless of whether the added layers are linear or MLP layers. This observation provides further evidence for the resemblance between DLNs and the deeper layers in MLPs. Additionally, it shows that DLNs can benefit from increased depth, similar to their counterparts in traditional nonlinear networks.
5.2 Experimental Verification of Assumption 3 and Theorem 4
@@
In all experiments, we fix the number of classes $K=3$ and the number of training samples $N=30$, with varying hidden dimensions $d\geq N$ as specified in each figure. To generate input data $\bm{X}\in\mathbb{R}^{d\times N}$ that satisfies Assumption 2, we first generate a matrix $\bm{A}\in\mathbb{R}^{d\times N}$ with entries i.i.d. sampled from the standard normal distribution and a matrix $\bm{B}\in\mathbb{R}^{d\times N}$ with entries i.i.d. sampled from the uniform distribution on $[0,1]$. Then, we apply a compact SVD to $\bm{A}$ and set ${\bm{X}}_{0}\in\mathbb{R}^{d\times N}$ as the left singular matrix. Next, we obtain a matrix $\bm{N}$ by normalizing $\bm{B}$ such that its Frobenius norm is 1. Finally, we generate $\bm{X}$ via $\bm{X}={\bm{X}}_{0}+\bm{N}$ such that $\bm{X}$ is nearly orthogonal satisfying Assumption 2.
@@
5.2.1 Experiments for Verifying Assumption 3
In this subsection, we conduct experiments to verify Assumption 3 and corroborate the discussions in Section 3.2 based on the above experimental setup. Specifically, we train DLNs using orthogonal initialization in (18) with $\xi=0.1$ for various network depths and widths. Given a DLN, we train it 10 times with different random seeds. In each run, we terminate training once the training loss is less than $10^{-11}$. We plot the following metrics over 10 runs against depth and width in Figure 7(a) and Figure 7(b), respectively:
[…]
$\text{variance of singular values of all weights:}\ \frac{1}{(L-1)(d-2K)}\sum_{l=1}^{L-1}\sum_{i=K+1}^{d-K}\left(\sigma_{i}(\bm{W}_{l})-\mu\right)^{2},$
where
$\mu=\frac{1}{(L-1)(d-2K)}\sum_{l=1}^{L-1}\sum_{i=K+1}^{d-K}\sigma_{i}(\bm{W}_{l}).$
It can be observed from Figure 7 that the weights of a DLN with different depths and widths trained under our settings approximately satisfy Assumption 3. This supports our discussions in Section 3.2.
5.2.2 Experiments for Verifying Theorem 4
@@
Now, we conduct experiments to validate Theorem 4. Specifically, we train 6-layer DLNs using GD with varying initialization scales $\xi=0.1,0.3,0.5$, respectively. Then, we plot the metrics of within-class compression $C_{l}$ and between-class discrimination $D_{l}$ against layer index in Figure 10(a) and Figure 10(b), respectively. It is observed that the feature compression metric $C_{l}$ decreases exponentially, while the feature discrimination metric $D_{l}$ increases linearly w.r.t. the layer index. In the top row of Figure 10, the solid blue line is plotted by fitting the values of $C_{l}$ at different layers, and the green and orange dash-dotted lines are plotted according to the lower and upper bounds in (15), respectively. We can observe that the solid blue line is tightly sandwiched between the green and orange dash-dotted lines. This indicates that (14) and (15) provide a valid and tight bound on the decay rate of the feature compression metric. According to the bottom row of Figure 10, we conclude that (16) provides a valid bound on the growth rate of the feature discrimination metric.
Moreover, we respectively train 6-layer DLN and 10-layer MLP networks with hidden dimension $d=50$ via GD with an initialization scale $\xi=0.3$. In Figure 5, we plot the feature compression and discrimination metrics on both DLN and MLP networks, respectively. We observe that the exponential decay of feature compression and the linear increase of feature discrimination hold exactly in DLNs and approximately in nonlinear networks.
@@ 5.3.1 Empirical Results Beyond Theory
  • Progressive feature compression and discrimination in nonlinear deep networks. We train 8-layer and 16-layer MLP networks, respectively. After training, we plot the metrics of feature compression and discrimination defined in (1) in Figure 11. It is observed that the feature compression metric $C_{l}$ decays at an approximate geometric rate in nonlinear networks, while the feature discrimination metric $D_{l}$ increases at an approximate linear rate. Additionally, it is worth mentioning that a similar “law of separation” phenomenon has been reported in he2023law; li2022principled on nonlinear networks, but their results are based upon a different metric of data separation.
@@
Beyond these findings, we conduct additional experiments on the SST-5 text dataset (socher2013recursive) and the ImageNet (deng2009imagenet) dataset to further support our claims. These results are reported in Section B.2.
(a) An 8-layer MLP.
(b) A 16-layer MLP.
@@
(a) A 5-layer DLN with Kaiming Init.
(b) An 8-layer DLN with Kaiming Init.
@@ 5.3.2 Implication on Transfer Learning
In this subsection, we experimentally substantiate our claims on transfer learning in Section 3 using practical nonlinear networks and real datasets. The results demonstrate that features before projection heads are less collapsed and exhibit better transferability. To support our argument, we conduct our experiments based on the following setup.
@@
We employ a ResNet18 backbone architecture (he2016deep), incorporating $t\in\{1,2,3,4,5\}$ layers of projection heads between the feature extractor and the final classifier, respectively. Here, one layer of the projection head consists of a linear layer followed by a ReLU activation layer. These projection layers are only used in the pre-training phase. On the downstream tasks, they are discarded, and a new linear classifier is trained on the downstream dataset.
@@
We use the CIFAR-100 and CIFAR-10 dataset in the pre-training and fine-tuning tasks, respectively. We train the networks using the Rescaled-MSE loss (hui2020evaluation), with hyperparameters set to $k=5$ and $M=20$ for 200 epochs. During pre-training, we employed the SGD optimizer with a momentum of 0.9, a weight decay of $5\times 10^{-4}$, and a dynamically adaptive learning rate ranging from $10^{-2}$ to $10^{-5}$, modulated by a CosineAnnealing learning rate scheduler (loshchilov2017sgdr). During the fine-tuning phase, we freeze all the parameters of the pre-trained model and only conduct linear probing. In other words, we only train a linear classifier on the downstream data for an additional 200 epochs. We run each experiment with 3 different random seeds.
@@
Figure 13: Adding projection heads mitigates feature collapse and improves transfer accuracy. We pre-train ResNet18 backbones with 1 to 5 layers of projection heads on the CIFAR-100 dataset. After pre-training, we drop the projection heads and compute the metric $C_{l}$ on the learned features. Finally, we linearly probe the pre-trained models on the CIFAR-10 dataset and plot the evolution of $C_{l}$ alongside the corresponding transfer learning performance against the number of projection head layers. The shaded area represents the standard deviation of 33 random seeds. The plot shows a clear trend that adding more layers of projection heads mitigates feature collapse and improves transfer learning results.
@@ 6 Conclusion
In this work, we studied hierarchical representations of deep networks by analyzing their intermediate features. In the context of training DLNs for solving multi-class classification problems, we examined how features evolve across layers through within-class feature compression and between-class feature discrimination. We showed that under mild assumptions on the input data and trained weights, each layer of a DLN progressively compresses within-class features at a geometric rate and discriminates between-class features at a linear rate w.r.t. the layer index. Moreover, we discussed the implications of our results for deep learning practices, including interpretation, architecture design, and transfer learning. Our extensive experimental results on synthetic and real data sets not only support our theoretical findings but also highlight their relevance to deep nonlinear networks.
@@
Our work opens several interesting directions for future work. First, our extensive experiments have demonstrated analogous patterns of progressive compression and discrimination phenomena in deep nonlinear networks. Extending our analysis to deep nonlinear networks emerges as a natural and promising direction for future research. Second, yu2023white; geshkovski2023emergence demonstrated that the outputs of self-attention layers in transformers exhibit a similar progressive compression phenomenon. It would be interesting to study this phenomenon in transformers based on our proposed framework.
Acknowledgments and Disclosure of Funding
@@
XL, CY, PW, and QQ acknowledge support from NSF CAREER CCF-2143904, NSF CCF-2212066, NSF CCF-2212326, NSF IIS 2312842, ONR N00014-22-1-2529, and an AWS AI Award. QQ also acknowledges the Google Research Scholar Award. PW and LB acknowledge support from DoE award DE-SC0022186, ARO YIP W911NF1910027, and NSF CAREER CCF-1845076. ZZ acknowledges support from NSF grants CCF-2240708, IIS-2312840, and IIS-2402952. WH acknowledges support from the Google Research Scholar Program. Results presented in this paper were obtained using CloudBank, which is supported by the NSF under Award #1925001. The authors acknowledge valuable discussions with Dr. Weijie Su (Upenn), Dr. John Wright (Columbia), Dr. Rene Vidal (Upenn), Dr. Hangfeng He (U. Rochester), Mr. Zekai Zhang (U. M), Dr. Yaodong Yu (UC Berkeley), and Dr. Yuexiang Zhai (UC Berkeley) at different stages of the work.
Appendix A Proofs of Main Results in Section 3